# Fluid dynamics, Statistical physics, Theoretical physics

### v1

@@ Abstract
Strongly nonlinear dynamics, from fluid turbulence to quantum chromodynamics, have long constituted some of the most challenging problems in theoretical physics. This review describes a unified theoretical framework, the loop space calculus, which offers an analytical approach to these problems. The central idea is a shift in perspective from pointwise fields to integrated loop observables, a transformation that recasts the governing nonlinear equations into a universal linear diffusion equation in the space of loops. This framework, supported by recent mathematical analysis, is analytically solvable and yields an exact, parameter-free solution for decaying hydrodynamic turbulence—the Euler ensemble—which is shown to be dual to a solvable string theory. The theory’s predictions include: (i) the unification of spatial and temporal scaling laws, which are shown to be governed by two related, infinite spectra of intermittency and decay exponents, respectively, both derived from the nontrivial zeros of the Riemann zeta function; (ii) a first-order phase transition in magnetohydrodynamic (MHD) turbulence; and (iii) the formation of quantized, concentric shells in passive scalar mixing. The theory also predicts log-periodic oscillations in correlation functions, effects not described by standard phenomenology, for which there is now emerging evidence from high-precision turbulence experiments. The appearance of identical mathematical structures as solutions to the turbulent regime of Yang-Mills gradient flow points towards the broad applicability of this approach. The framework also yields a new type of minimal surface that solves the QCD loop equations in the limit of large loops.
@@ A Brief Historical Context
This review is organized as follows. The discussion begins with an introduction to the general loop space approach, applicable to both abelian fields in fluid dynamics and nonabelian fields in Yang-Mills theory. The approach is then applied to a hierarchy of turbulence problems, beginning with the foundational case of decaying hydrodynamic turbulence, then the transport of passive scalars, and finally the more complex case of magnetohydrodynamic (MHD) turbulence. A subsequent section situates these analytical solutions within the field’s historical context and clarifies their relationship to classical paradigms of turbulence theory. The review concludes by examining the appearance of the same mathematical structures as exact solutions in fundamental gauge theory and exploring the framework’s broad applicability.
@@ 2 The Loop Space Approach: From Nonlinear PDEs to Linear Diffusion
The transformation of both equations to the loop space diffusion equation is based on a fundamental mathematical identity, known as the Magnus expansion for path-ordered exponentials [Magnus(1954)]. As applied to our case of the loop functionals for fluid dynamics and the gauge theory, it allows us to write both of them as the Magnus ordered exponential times the unit operator
@@
This geometric formula, proven in the Appendix, is identical in abelian and non-abelian cases, except that there is a trace taken in the Yang-Mills case. In both cases, it reveals the geometric meaning of the loop functional as a parallel transport operator around a loop $C$, reducing to a c-number for a closed loop.
@@ 4 Application I: Decaying Hydrodynamic Turbulence
Figure 2: Verification of the Euler-ensemble theory against new high-resolution 4K DNS data. The theory’s predictions are compared with recent high-precision Direct Numerical Simulations (DNS) of decaying isotropic turbulence [Sreenivasan and Rodhiya, Akash(2026)]. (Top Left): The predicted parabolic relation between time $t$ and the integral length scale $L_{M}(t)$. The simulation data (blue line) is in excellent agreement with a parabolic fit (orange dashed line), confirming the theoretical $L_{M}\propto t^{1/2}$ law. (Top Right): The parameter-free prediction for the effective index $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$ (orange line). The theory captures the universal curvature seen in the DNS data (blue line, with uncertainty band), a feature not described by classical K41 scaling. (Bottom Left & Bottom Right): The parameter-free prediction for the energy decay, $E\propto L_{M}^{-5/2}$ (gray dash-dotted line). This prediction is tested against DNS data at two different Reynolds numbers: $Re_{\lambda}=145$ (Bottom Left) and $Re_{\lambda}=93$ (Bottom Right). The simulations show $E\propto L_{M}^{-2.36}$ and $E\propto L_{M}^{-2.40}$ respectively, in close agreement with the theoretical exponent of $2.5$.
@@ 4.1 Key Prediction 1: The Energy Spectrum and Intermittency
Furthermore, the log-log slope of $E(t)$ versus $L(t)$ in the DNS data (showing exponents of 2.36-2.36 and 2.40-2.40 in Fig. 2) deviates slightly from the leading theoretical exponent of 5/2-5/2. The theory also explains this deviation, as the leading power law is just the first "ground state" term in an infinite series. The next correction, corresponding to the $L^{-11/2}$ exponent in the theoretical spectrum [Migdal(2024h)], accounts for this deviation. This subleading term plays a larger role at smaller $L_{M}$ (at the higher-energy $E\sim 10^{-2}$ end of the plot), leading to a larger deviation. As $L_{M}$ grows (at the lower energy $E\sim 10^{-4}$ end), the correction becomes less significant, and the DNS slope approaches the pure 5/2-5/2 law. When higher-resolution data becomes available, it should be compared against the full theoretical curve, which accounts for all corrections by numerically evaluating the Mellin-Barnes integral. As shown in [Migdal(2024h)] using preliminary DNS data, this full solution provides a perfect match across all scales.
@@ 4.2 Key Prediction 2: The Geometry of Mixing
The Euler ensemble also dictates the behavior of a passive scalar (e.g., temperature or a dye) advected by the turbulent flow [Migdal(2025d)]. For a localized initial condition, the scalar does not spread into a simple Gaussian cloud. We derived the loop equation for the evolution of the passive scalar $T(r,t)$ from conventional diffusion-advection equations, coupled with the loop equations for the velocity circulation. The resulting solution for the scalar density in the extreme turbulent limit is quite unexpected. Instead of a smooth Gaussian cloud, it forms a series of expanding, quantized concentric shells (see Figure 3. The radial profile of the scalar concentration is piecewise-parabolic, and the shell structure is organized by the Euler totient function, which counts the number of valid star polygons of a given complexity. This sharp, geometric prediction offers a clear target for future experiments.
@@ 5 Application II: Magnetohydrodynamic (MHD) Turbulence
The loop space solution can be generalized to more complex systems, such as magnetohydrodynamics (MHD), which describes the dynamics of conducting fluids, such as astrophysical plasmas [Migdal(2025c)]. The system is now described by two coupled loop functionals, one for the velocity circulation ($\Gamma_{v}$) and one for the magnetic vector potential circulation ($\Gamma_{a}$), each governed by an interacting Euler ensemble.
@@ 6 The Yang-Mills Connection and Universal Duality
The final step reveals the framework’s profound universality. The Yang-Mills gradient flow, a central tool for understanding the vacuum structure and confinement problem in QCD, is governed by a non-Abelian version of the same loop dynamics [Migdal(2025b)]. Applying our framework, we find exact solutions for the evolution of the Wilson loop.
Remarkably, one of the two solutions mirrors those found in 3-dimensional turbulence. At the same time, another one– a Hodge-dual minimal area– is present only in the non-Abelian Yang-Mills theory on four dimensions. The YM loop equation in four dimensions admits:
  • A stationary, stable fixed point corresponds to an area law $W[C]\propto\exp\left(-\kappa(A_{+}[C]+A_{-}[C])\right)$ with the area of self-dual and anti-dual minimal surfaces $A_{\pm}[C]$ in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ space, bounded by $C$.
    1. 1.
      This is a new four-dimensional geometric object, analogous to the gauge theory instanton solution (see [Migdal(2025b)] and Appendix of this review).
    2. 2.
      For a planar loop, the area of this self-dual surface is $2\sqrt{2}$ times larger than the geometric area inside the loop.
    3. 3.
      For a general loop, it grows proportionally to the ordinary minimal area inside the 4-dimensional loop, providing a quark confining mechanism.
  • A universal decaying trajectory that is, once again, the Euler ensemble.
    1. 1.
      The spontaneous quantization of parameters by rational numbers describes the statistics of the Wilson loop.
    2. 2.
      The Wilson loop is expressed in terms of the average over the Euler ensemble of solutions to the universal ODE on a circle, with parameters depending upon the distribution of the rational numbers, measured by Euler totients.
@@ 8 A New Perspective on Classical Models in Decaying Turbulence
The Hodge-Dual Surface as a Theoretical Area law in QCD
Why would we need another area law if the existing lattice QCD simulations confirm the conventional area law? This is a crucial question that touches upon the relationship between theoretical solutions and numerical validation. There are two fundamental reasons.
  • The Principle of First-Principles Theory: As a matter of principle, the goal of theoretical physics is to find analytical solutions to fundamental equations. Numerical simulations, while indispensable, serve to validate or falsify proposed theories; they cannot discover new laws of nature but rather adjudicate the theoretical conjectures. While a simple area law provides a good fit to current data, it does not solve the loop equations. The Hodge-dual minimal surface, by contrast, is derived as an exact solution. It provides the theoretical "why" behind the observed "what," a distinction that is central to fundamental understanding. In the same way that the Euler ensemble provided a microscopic theory that could replace phenomenological models in turbulence, an exact string solution is needed to provide a definitive foundation for non-perturbative QCD.
  • The Limitations of Current Numerical Experiments: There are strong reasons to believe that current lattice simulations, despite their power, are not yet in a regime to observe the key predictions that distinguish the Hodge-dual surface from a simple area law.
    • The Challenge of Scale: The largest lattice simulations in QCD are performed on grids with fewer than 100 points per dimension. By comparison, turbulence DNS now reaches resolutions of $32,000^{3}$. Even at that immense scale, it was difficult to definitively rule out flexible phenomenological models such as multifractal scaling. Only a direct comparison with the parameter-free predictions of the Euler ensemble made the discrepancies apparent. It is plausible that the subtle, shape-dependent effects of the true confining string are simply beyond the reach of current lattice resolutions.
    • The Challenge of Geometry: Crucially, the Wilson loops that have been measured in lattice experiments are overwhelmingly planar. For a planar loop, the area of the Hodge-dual surface is $2\sqrt{2}$ times the enclosed Euclidean area. This constant prefactor is indistinguishable from an adjustment to the string tension and cannot be used to falsify the model. The definitive test would be to measure the Wilson loop for a non-planar contour, such as a "twisted rectangle" in which one side is rotated by 90 degrees relative to the original plane. For such a screw-like loop, the Hodge dual and the Euclidean minimal surface predict different areas. While such a measurement may be computationally prohibitive today, it remains the clear and necessary experiment to test the theory’s geometric predictions.
@@ 9 Conclusion and Outlook
Among these avenues, the theory’s formulation of the confining string is of particular interest for quantum chromodynamics. The framework yields an exact solution to the QCD loop equations for non-intersecting loops in the form of a Hodge-dual minimal surface [Migdal(2025b)]. This geometric object, by its very construction, satisfies the loop equations in four-dimensional space, unlike the conventional Euclidean minimal surface, which violates the full set of Yang-Mills loop equations. Our Hodge-dual minimal surface provides a more nuanced physical picture. While both surfaces yield a confining area law, the Hodge-dual area shows a non-trivial dependence on the shape of non-planar Wilson loops. This shape dependence provides a first-principles mechanism for calculable, shape-dependent corrections to the linear confining potential, moving beyond idealized models toward a more complete description of the quark-quark force.
@@ Appendix A Essentials of the loop space calculus
The full derivations and technical details can be found in our recent paper [Migdal(2025a)].
@@
The area derivative is defined by the non-abelian Stokes theorem. For a detailed derivation, see [Migdal(2025a)], Sec. 3.3.
@@
Operator representation of the Wilson loop
The Wilson loop can be represented as a trace of a path-ordered exponential of covariant derivative operators. For details, see [Migdal(2025a)], Sec. 3.4, Eqs. (12-15).
$C_{\mu}(\theta)$ $=x_{\mu}+\int_{0}^{\theta}d\theta^{\prime}\dot{C}_{\mu}(\theta^{\prime});\quad\int_{0}^{2\pi}d\theta^{\prime}\dot{C}_{\mu}(\theta^{\prime})=0;$ (58)
$W[C]\otimes\mathbb{I}$ $=\frac{1}{N}\left\langle\hbox{tr}\,P\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)D_{\mu}(x)\right)\right\rangle;$ (59)
$D_{\mu}(x)$ $=\partial_{x_{\mu}}+A_{\mu}(x)$ (60)
The "dot derivative" $\frac{\delta}{\delta\dot{C}_{\mu}(t\pm)}$ brings down a $D_{\mu}$ operator. For details, see [Migdal(2025a)], Sec. 3.4, Eq. (20).
$\left.P\exp\left(\mathrm{i}\int_{t}^{t+2\pi}ds\dot{C}_{\mu}(\theta)D_{\mu}(C(t))\right)\right\rangle$ (61)

### v2

@@ Abstract
Strongly nonlinear dynamics, from fluid turbulence to quantum chromodynamics, have long constituted some of the most challenging problems in theoretical physics. This review describes a unified theoretical framework, the loop space calculus, which offers an analytical approach to these problems. The central idea is a shift in perspective from pointwise fields to integrated loop observables, a transformation that recasts the governing nonlinear equations into a universal linear diffusion equation in the space of loops. This framework, supported by recent mathematical analysis, is analytically solvable and yields an exact, parameter-free solution for decaying hydrodynamic turbulence—the Euler ensemble—which is shown to be dual to a solvable string theory. The theory’s predictions include: (i) the unification of spatial and temporal scaling laws, governed by two related, infinite spectra of intermittency and decay exponents derived from the nontrivial zeros of the Riemann zeta function; (ii) a first-order phase transition in magnetohydrodynamic (MHD) turbulence; and (iii) the formation of quantized, concentric shells in passive scalar mixing. The theory also predicts log-periodic oscillations in correlation functions — effects not captured by standard phenomenology — for which there is now emerging evidence from high-precision turbulence experiments. The appearance of identical mathematical structures as solutions to the turbulent regime of Yang–Mills gradient flow points to the broad applicability of this approach. The framework also yields a new type of analytic Hodge-dual matrix surface that solves the Yang–Mills fixed-point loop equation by harmonic map, opening the way for a geometric formulation of QCD string theory.
@@ A Brief Historical Context
This review is organized as follows. The discussion begins with an introduction to the general loop space approach, applicable to both abelian fields in fluid dynamics and nonabelian fields in Yang-Mills theory. The approach is then applied to a hierarchy of turbulence problems, beginning with the foundational case of decaying hydrodynamic turbulence, then the transport of passive scalars, and finally the more complex case of magnetohydrodynamic (MHD) turbulence. A subsequent section situates these analytical solutions within the field’s historical context and clarifies their relationship to classical paradigms of turbulence theory. A particularly striking consequence of this framework emerges in gauge theory. In the Yang–Mills gradient flow, the same loop-space formalism leads to an exact analytic solution—the Hodge-dual matrix surface—which solves the fixed-point Yang–Mills loop equation by harmonic map [Migdal(2025a)]. This result establishes a concrete geometric foundation for the QCD confining string and provides a direct analytical bridge between turbulence, gauge theory, and string dualities.
@@ 2 The Loop Space Approach: From Nonlinear PDEs to Linear Diffusion
Operator identity transforms nonlinear flow into loop space diffusion
The transformation of both equations to the loop space diffusion equation is based on the operator calculus invented by Feynman [Feynman(1951)]. As applied to our case of the loop functionals for fluid dynamics and the gauge theory, it allows us to write both of them as the path ordered exponential times the unit operator
@@
This fundamental geometric identity, proven in the [Migdal(2025b)] and our Appendix A, is identical in abelian and non-abelian cases, except that a group space trace is taken in the Yang-Mills case. In both cases, it reveals the geometric meaning of the loop functional as a parallel transport of the covariant derivative operator around a loop $C$, reducing to a c-number for a closed loop. The reduction of the ordered exponential to a c-number is also supported by the Magnus expansion [Magnus(1954)]. This distinction is subtle but critical. In much of the mathematical literature such as [Göckeler and Schücker(1987)], the Wilson loop (the L.H.S. of our identity) is taken as the definition of parallel transport. Our proof, however, establishes this correspondence as a rigorous equality derived from the more fundamental, anchored operator (the R.H.S.). This approach mirrors a historical pattern in theoretical physics where a concept, once viewed as a powerful analogy, is later shown to be a provable, first-principles identity.
@@ 4 Application I: Decaying Hydrodynamic Turbulence
Figure 2: Verification of the Euler-ensemble theory against new high-resolution 4K DNS data [Sreenivasan and Rodhiya, Akash(2026)]. (Top Left): The predicted parabolic relation between time $t$ and the integral length scale $L_{M}(t)$. The simulation data is in excellent agreement with a parabolic fit. (Top Right): The parameter-free prediction for the effective index $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$. The theory captures the universal curvature seen in the DNS data. (Bottom): Verification of the full energy decay law at $Re_{\lambda}=145$. The DNS data (blue line) deviates from the leading-order term of the theory ($E\propto L_{M}^{-5/2}$, gray dash-dotted line). However, it is in excellent agreement with the complete theoretical solution (orange line), which includes all sub-leading exponents from the Mellin-Barnes integral. Only the two unknown scales (energy and length) were fitted to match the DNS data with the theory; no dimensionless parameters were changed. Such a perfect fit with only two parameters for about a thousand data points provides a high level of confidence.
@@ 4.1 Key Prediction 1: The Energy Spectrum and Intermittency
Furthermore, the log-log slope of $E(t)$ versus $L(t)$ in the DNS data shows a slight, systematic deviation from the leading-order theoretical exponent of 5/2-5/2. This deviation is not a flaw; it is perfectly explained by the theory.
As shown in Figure 2 (Bottom), the leading-order term (gray dash-dotted line) acts as an asymptote, while the complete theoretical curve (orange line)—which includes all sub-leading exponents from the numerically evaluated Mellin-Barnes integral—provides a precise match to the $Re_{\lambda}=145$ DNS data over the entire range. This fit, which matches thousands of data points using only two physical scale parameters (for overall energy and length), provides a high level of confidence in the solution. It visually confirms that the observed energy decay law is composed of the "ground state" term ($L_{M}^{-5/2}$) plus a hierarchy of calculable, quantum-like corrections, just as the theory predicts.
Presumably, the minor deviations from the theoretical curve at the far left end ($\log L_{M}<3.2,\log E>2$) are explained by the early stages of energy decay, when statistical equilibrium has not yet been achieved. The initial K41 spectrum would have contaminated the data, leading to deviations from the turbulent attractor.
@@ 4.2 Key Prediction 2: The Geometry of Mixing
The Euler ensemble also dictates the behavior of a passive scalar (e.g., temperature or a dye) advected by the turbulent flow [Migdal(2025e)]. For a localized initial condition, the scalar does not spread into a simple Gaussian cloud. We derived the loop equation for the evolution of the passive scalar $T(r,t)$ from conventional diffusion-advection equations, coupled with the loop equations for the velocity circulation. The resulting solution for the scalar density in the extreme turbulent limit is quite unexpected. Instead of a smooth Gaussian cloud, it forms a series of expanding, quantized concentric shells (see Figure 3. The radial profile of the scalar concentration is piecewise-parabolic, and the shell structure is organized by the Euler totient function, which counts the number of valid star polygons of a given complexity. This sharp, geometric prediction offers a clear target for future experiments.
@@ 5 Application II: Magnetohydrodynamic (MHD) Turbulence
The loop space solution can be generalized to more complex systems, such as magnetohydrodynamics (MHD), which describes the dynamics of conducting fluids, such as astrophysical plasmas [Migdal(2025d)]. The system is now described by two coupled loop functionals, one for the velocity circulation ($\Gamma_{v}$) and one for the magnetic vector potential circulation ($\Gamma_{a}$), each governed by an interacting Euler ensemble.
@@ 6 The Yang-Mills Connection and Universal Duality
The final step reveals the framework’s profound universality. The Yang-Mills gradient flow, a central tool for understanding the vacuum structure and confinement problem in QCD, is governed by a non-Abelian version of the same loop dynamics [Migdal(2025c)]. Applying our framework, we find exact solutions for the evolution of the Wilson loop.
In the Yang–Mills gradient flow, the loop-space calculus leads not only to the diffusion equation for the Wilson loop but also to its stationary fixed point, which defines the confining string. The analytic solution is provided by the Hodge-dual matrix surface, obtained by the harmonic-map method. This surface satisfies the self-duality relation
$\Sigma_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta},$ (48)
and minimizes the dual area functional under this constraint. Its area $S[C]$ exactly solves the non-perturbative fixed-point Yang–Mills loop equation,
$\partial_{\mu}\frac{\delta S[C]}{\delta\sigma_{\mu\nu}(\theta)}=0,$ (49)
ensuring that $\exp\left(-\kappa S[C]\right)$ is annihilated by the loop operator. This construction yields a parity-even, factorized form of the Wilson functional,
$W[C]=\exp\left(-\kappa(S_{+}[C]+S_{-}[C])\right),$ (50)
where $S_{\pm}[C]$ are the areas of the analytic self-dual and anti-self-dual surfaces bounded by the same loop.
For planar loops, the dual area is $|S|=2\sqrt{2}|D|$, reproducing the confining area law, while for general nonplanar or twisted loops, the shape dependence distinguishes it from the Euclidean minimal surface. The harmonic-map solution [Migdal(2025a)] clarifies the geometry: the Hodge-dual surface corresponds to a harmonic embedding $X:\mathcal{D}\!\to\!M_{N}(\mathbb{C})$ with the induced conformal metric $g_{ab}=e^{\varphi}\delta_{ab}$, where the Liouville field $\varphi$ governs local scale fluctuations. This analytic result provides the first explicit geometric realization of the Yang–Mills confining string and establishes a bridge between gauge theory, dual string geometry, and the loop-space formulation of turbulence.
@@ 8 A New Perspective on Classical Models in Decaying Turbulence
The Hodge–Dual Surface as a Theoretical Area Law in QCD
Why introduce a new area law when lattice QCD already confirms the conventional one? The answer lies in distinguishing a numerical observation from an analytic solution. Two arguments are central.
  • Analytic foundation. Numerical simulations can test or falsify models but cannot derive the law itself from first principles. The analytic Hodge–dual matrix surface, constructed by the harmonic–map method, exactly solves the fixed–point Yang–Mills loop equation [Migdal(2025a)]. It provides the theoretical explanation for the confining area law rather than an empirical fit, offering a genuine analytic counterpart to the numerical area law of lattice QCD.
  • Beyond current lattice reach. Present lattice resolutions are insufficient to test the distinctive, shape–dependent predictions of the Hodge–dual surface. Existing calculations mostly involve planar Wilson loops, for which the dual area is $2\sqrt{2}$ times the Euclidean area—a factor that can be absorbed into the string tension. The crucial test would involve nonplanar, “twisted” contours where the dual and Euclidean minimal surfaces yield different areas. Such computations remain demanding but would provide a definitive check of the analytic theory once feasible.
In this sense, the Hodge–dual matrix surface does not replace the lattice area law but explains its origin and predicts measurable deviations for nonplanar geometries, establishing an analytic framework for the QCD confining string. Being an exact analytic solution, it also provides a natural starting point for quantizing the string, beginning from this holomorphic minimal-surface solution. The resulting quantized theory offers a systematic path to computing the hadron spectrum from first principles, extending far beyond the phenomenology of the linearly rising potential.
@@ 9 Conclusion and Outlook
Among these avenues, the theory’s formulation of the confining string is of particular interest for quantum chromodynamics. The framework yields an exact solution to the QCD loop equations for non-intersecting loops in the form of a Hodge-dual minimal surface [Migdal(2025c)]. This geometric object, by its very construction, satisfies the loop equations in four-dimensional space, unlike the conventional Euclidean minimal surface, which violates the full set of Yang-Mills loop equations.
The analytic Hodge–dual matrix surface derived by the harmonic–map method [Migdal(2025a)] provides the geometric core of the Yang–Mills fixed–point solution, completing the dual correspondence between turbulence and confinement. Beyond this analytic construction, the holomorphic reduction yields an explicit Confinement Theorem. Define the spectral ratio
$r[C]$ $=\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}\in\Big[\tfrac{1}{2},\,1\Big],$ (60)
where $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\lambda_{4}$ are the ordered eigenvalues of the Douglas–Gram tensor
$G_{\mu\nu}[C]$ $=\sum_{n>0}n\,\hat{C}_{n,\mu}\,\hat{C}_{n,\nu},\qquad\hat{C}_{n,\mu}=\tfrac{1}{2\pi}\!\int_{0}^{2\pi}\!C_{\mu}(\theta)e^{-in\theta}\,d\theta.$ (61)
From the analytic solution, the Hodge–dual minimal area of the surface and the Euclidean minimal area of the same boundary loop are expressed as
$S_{\chi}[C]$ $=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ $S_{\mathrm{E}}[C]$ $=\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4},$ (62)
so that their ratio directly yields the spectral ratio $r[C]$. Hence the effective string tension,
$\sigma_{\mathrm{eff}}[C]$ $=\frac{S_{\chi}[C]}{S_{\mathrm{E}}[C]}=2\sqrt{2}\,r[C],$ (63)
obeys the universal bounds
$\sqrt{2}\ \leq\ \sigma_{\mathrm{eff}}[C]\ \leq\ 2\sqrt{2},$ (64)
for all loops. The numerical value of the lower bound is simple, but its existence—independent of loop size and shape—is highly nontrivial and constitutes the confinement statement in this framework. All residual shape dependence of the string tension is encoded by the spectral ratio $r[C]$: planar loops (rank–2 $G$) saturate the upper bound $r[C]=1$, while isotropic spectra saturate the lower bound $r[C]=\tfrac{1}{2}$. This explicit formula for the shape dependence of the quark string tension opens the way to computing the rising Regge trajectories in QCD from first principles.
While the same mathematical formalism connects turbulence, gauge theory, and string duality, these correspondences should not be interpreted as a unification of the underlying physical systems but rather as a shared geometric language linking universal features of their non-linear dynamics.
@@ Appendix A Essentials of the loop space calculus
The full derivations and technical details can be found in our recent paper [Migdal(2025b)].
@@
The area derivative is defined by the non-abelian Stokes theorem. For a detailed derivation, see [Migdal(2025b)], Sec. 3.3.
@@
A.1 Operator representation of the Wilson loop
The loop functional can be represented as a trace of a path-ordered exponential of covariant derivative operator. This subsection proves the fundamental operator identity that enables this. The full derivation, originally presented in [Migdal(2025b)] (Sec. 3.4), is reproduced here with all necessary mathematical details. We also fix some typos in that paper.
Theorem 1 (Operator-Holonomy Identity).
Let $D_{\mu}(x)=\mathbb{I}_{G}\partial_{\mu}+A_{\mu}(x)$ be the covariant derivative at an arbitrary base point $x\in\mathbb{R}^{d}$, where $\partial_{\mu}$ is the partial derivative in Hilbert space and $A_{\mu}(x)$ is a matrix-valued operator in the group space $G$, and $\mathbb{I}_{G}$ is a unity matrix in that group space. Let $C$ be a smooth, closed loop parameterized by $\theta\in[0,2\pi]$ with $\dot{C}_{\mu}(\theta)$ as its velocity, such that $C(\theta)=x+\int_{0}^{\theta}d\theta^{\prime}\dot{C}_{\mu}(\theta^{\prime})$, and $\int_{0}^{2\pi}d\theta^{\prime}\dot{C}_{\mu}(\theta^{\prime})=0$.
The path-ordered exponential of the covariant derivative operator $D_{\mu}(x)$ along this loop is equal to the Wilson loop (holonomy) along the path $C(\theta)$ multiplied by the identity operator $\mathbb{I}$ in the Hilbert space.
$\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)D_{\mu}(x)\right)=\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)A_{\mu}(C(\theta))\right)\otimes\mathbb{I}$ (66)
Proof.
The proof relies on discretizing the path and using the "disentangling" identity, which is a consequence of Feynman’s operator calculus [Feynman(1951)].
1. Discretization
The path-ordered exponential on the L.H.S. is formally defined as the limit of a product integral:
$\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)D_{\mu}(x)\right)=\lim_{N\to\infty}\prod_{k=N\to 1}\exp\left(\Delta\theta_{k}\dot{C}_{\mu}(\theta_{k})D_{\mu}(x)\right)$ (67)
where $\Delta\theta_{k}\to 0$ and the product is ordered from right to left (i.e., $k=1$ is the rightmost operator). Let $dC_{\mu}(\theta)=d\theta\dot{C}_{\mu}(\theta)$.
2. The Infinitesimal Disentangling
We analyze a single infinitesimal factor $\exp\left(dC_{\mu}D_{\mu}(x)\right)$. Using the Lie-Trotter product formula, $e^{A+B}=e^{A}e^{B}+O([A,B])$, we can split the operator:
$\exp\left(dC_{\mu}(\partial_{\mu}+A_{\mu}(x))\right)=\exp\left(dC_{\mu}\partial_{\mu}\right)\exp\left(dC_{\mu}A_{\mu}(x)\right)+O(d\theta^{2})$ (68)
The $O(d\theta^{2})$ terms vanish in the $N\to\infty$ limit. The $\exp\left(dC_{\mu}\partial_{\mu}\right)$ operator is an infinitesimal translation operator, $T_{dC}$. We now use the fundamental operator identity that defines a translation:
$T_{dC}\,f(x)=f(x+dC)\,T_{dC}$ (69)
Applying this to our $A_{\mu}(x)$ operator, we find the "hopping" identity:
$\exp\left(dC_{\mu}\partial_{\mu}\right)\exp\left(dC_{\mu}A_{\mu}(x)\right)=\exp\left(dC_{\mu}A_{\mu}(x+dC)\right)\exp\left(dC_{\mu}\partial_{\mu}\right)$ (70)
This is the "disentangling" from (A7) in the text. It allows us to move the (Abelian) derivative operator to the right, at the cost of shifting the argument of the (non-Abelian) $A_{\mu}$ operator.
3. Iteration and Re-ordering
We now apply this identity iteratively to the full product. Let $C_{k}=C(\theta_{k})=x+\int_{0}^{\theta_{k}}d\theta^{\prime}\dot{C}(\theta^{\prime})$. By periodicity, $C_{N}=x$.
$\prod_{k=N\to 1}\exp\left(dC_{k}\cdot D(x)\right)\approx\prod_{k=N\to 1}\left(\exp\left(dC_{k}\cdot A(C_{k})\right)\exp\left(dC_{k}\cdot\partial_{x}\right)\right)$ (71)
$=\left(e^{dC_{N}\cdot A(C_{N})}e^{dC_{N}\cdot\partial_{x}}\right)\left(e^{dC_{N-1}\cdot A(C_{N})}e^{dC_{N-1}\cdot\partial_{x}}\right)\dots\left(e^{dC_{1}\cdot A(C_{N})}e^{dC_{1}\cdot\partial_{x}}\right)$ (72)
We repeatedly "hop" all the $\exp\left(dC_{k}\cdot\partial_{x}\right)$ terms to the far right. Each time a $\partial_{x}$ operator passes an $A(C_{N})$ operator, it shifts its argument. The crucial point is that all $\partial_{x}$ operators are Abelian and commute with each other, while all $A(C_{k})$ operators are non-Abelian and remain path-ordered. Furthermore, the resulting argument $x_{k}$ of each $A$ in the product at kthk-th place will be
$x_{k}=C_{N}+\sum_{n=N\to k}dC_{n}=C_{k}$ (73)
This process separates the product into two distinct, ordered parts:
$=\left(\prod_{k=N\to 1}e^{dC_{k}\cdot A(C_{k})}\right)\left(\prod_{k=N\to 1}e^{dC_{k}\cdot\partial_{x}}\right)$ (74)
Taking the $N\to\infty$ limit, this becomes:
$=\left(\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)A_{\mu}(C(\theta))\right)\right)\left(\exp\left(\int_{0}^{2\pi}dC_{\mu}\partial_{\mu}\right)\right)$ (75)
4. The Closed Loop Condition
The first term is, by definition, the Wilson loop $W[C]$ (the holonomy), which is a c-number matrix. The second term is the identity operator in Hilbert space, because the loop $C$ is closed:
$\int_{0}^{2\pi}dC_{\mu}=\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)=C_{\mu}(2\pi)-C_{\mu}(0)=0$ (76)
Therefore, the operator becomes:
$\exp\left(\left(\int_{0}^{2\pi}dC_{\mu}(\theta)\right)\partial_{x_{\mu}}\right)=\exp\left(0\cdot\partial_{x_{\mu}}\right)=\mathbb{I}$ (77)
This completes the proof. We find that the Hilbert space operator on the L.H.S. is equal to the group space matrix (the Wilson loop) multiplied by the identity operator $\mathbb{I}$. ∎